An unfinished quilt in the shape of a triangle has an area of 27m2. If the base of the triangular quilt is 3m more than the height, what are the dimensions of the triangular quilt?
Answer by rothauserc(4718)

let h be the triangles height, then
:
h+3 is the base
:
(1/2) * (h+3) * h = 27
:
(h+3) * h = 54
:
h^2 +3h -54 = 0
:
(h+9) * (h-6) = 0
:
h=6, -9
:
we reject the negative value for h, then
:
the dimensions of the triangular quilt are height=6m, base=9m
: